### v2

# LLM-Coordination: Evaluating and Analyzing Multi-agent Coordination Abilities in Large Language Models

Abstract
The emergent reasoning and Theory of Mind (ToM) abilities demonstrated by Large Language Models (LLMs) make them promising candidates for developing coordination agents. In this study, we introduce a new LLM-Coordination Benchmark aimed at a detailed analysis of LLMs within the context of Pure Coordination Games, where participating agents need to cooperate for the most gain. This benchmark evaluates LLMs through two distinct tasks: (1) Agentic Coordination, where LLMs act as proactive participants for cooperation in 4 pure coordination games; (2) Coordination Question Answering (QA), where LLMs are prompted to answer 198 multiple-choice questions from the 4 games for evaluation of three key reasoning abilities: Environment Comprehension, ToM Reasoning, and Joint Planning. Furthermore, to enable LLMs for multi-agent coordination, we introduce a Cognitive Architecture for Coordination (CAC) framework that can easily integrate different LLMs as plug-and-play modules for pure coordination games. Our findings indicate that LLM agents equipped with GPT-4-turbo achieve comparable performance to state-of-the-art reinforcement learning methods in games that require commonsense actions based on the environment. Besides, zero-shot coordination experiments reveal that, unlike RL methods, LLM agents are robust to new unseen partners. However, results on Coordination QA show a large room for improvement in the Theory of Mind reasoning and joint planning abilities of LLMs. The analysis also sheds light on how the ability of LLMs to understand their environment and their partner’s beliefs and intentions plays a part in their ability to plan for coordination. Our code is available at https://github.com/eric-ai-lab/llm_coordination.
1 Introduction
In a wide range of activities, from daily tasks such as cooking to critical operations like rescue efforts, cooperation without mixed intentions is essential. These scenarios are examples of Pure Coordination Games, where all involved parties benefit from choosing strategies that are perfectly aligned, avoiding any conflict of interest. These games require agents to reason about their environment and plan while considering the beliefs and intentions of their partners. Recently, Large Language Models (LLMs) have demonstrated emergent planning abilities in both physical and virtual settings (Raman et al., 2022; Wang et al., 2023; Wu et al., 2023), impressive reasoning capabilities (Wei et al., 2022), and the hints of a Theory of Mind (Kosinski, 2023) making them promising candidates for developing coordination agents. Previous works have explored the use of LLMs for developing collaborative agents, yet the requisite conditions, strengths, and limitations of LLMs in coordination games remain unclear. In this study, we intend to bridge the gap by performing a comprehensive evaluation and analysis of the multi-agent coordination abilities of LLMs.
Therefore, we introduce a new LLM-Coordination Benchmark featuring two task settings for pure coordination games: 1. Agentic Coordination and 2. CoordinationQA. In Agentic Coordination, LLMs are scaffolded with components that allow them to act within actual game environments, providing a holistic evaluation of the competencies of LLMs to act as coordination agents. In CoordinationQA, LLMs have to answer a curated set of questions about edge-case scenarios drawn from coordination games where agents are required actively cooperate with their partners. The benchmark includes four collaborative games, providing a comprehensive analysis platform.
To enable LLMs for multi-agent coordination, we present a Cognitive Architecture for Coordination (CAC) framework that facilitates LLM interaction with game environments in a plug-and-play approach. CAC translates game elements into textual formats and leverages auxiliary LLMs for improved coordination to enable effective multi-agent collaboration. Our Experiments on the Agentic Coordination task with CAC reveal that Large Language Models are capable of understanding the game objectives, generating coherent reasoning for their next actions, and coordinating with partners across all coordination games. They exhibit these competencies without any training, fine-tuning, or few-shot examples.
A comparative analysis on the Agentic Coordination task reveals that LLM agents outperform Multi-agent RL solutions in games that require minimum theory-of-mind reasoning and focus on taking commonsense actions based on the environment. However, they struggle in more complex settings where agents need to actively consider their partner’s beliefs and intentions. We also observe that LLM agents are capable of collaborating with new partners, unlike self-play MARL methods (Carroll et al., 2019a; Bard et al., 2020) that fail to adapt to unseen agents.
For a more nuanced analysis of the coordination abilities of LLMs, we create the CoordinationQA Suite. This suite is designed to dissect the capabilities of LLMs in single-turn reasoning within coordination games, focusing on three key areas: Joint Planning, Theory of Mind (ToM), and Environment Comprehension. Joint Planning evaluates LLMs’ decision-making for optimal long-term outcomes, ToM questions probe the understanding of partner agents’ intentions and needs, and Environment Comprehension assesses knowledge of game rules and dynamics. Our findings on CoordinationQA show a marked performance gap between GPT-4-turbo and other LLMs across three question types. LLMs are most proficient in Environment Comprehension, indicating they understand game rules and states well. However, they face significant challenges in Theory of Mind Reasoning, with difficulty inferring others’ intentions and needs. This issue worsens in Joint Planning, where most LLMs underperform, some even worse than random choices. These results highlight LLMs’ limited reliability and effectiveness as coordination partners. Additionally, our analysis shows a moderate to strong correlation between Theory of Mind Reasoning, Environment Comprehension, and Joint Planning performance, underscoring their role in effective coordination.
In summary, our contributions are four-fold:
1.
We introduce the LLM-Coordination Benchmark for evaluating and analyzing LLMs in Pure Coordination Games, covering multi-turn Agentic Coordination and single-turn Coordination QA tasks.
2.
We develop the plug-and-play Cognitive Architecture for Coordination Framework to enable LLMs to effectively participate in complex, partially observable coordination games like Hanabi, demonstrating the first zero-shot application for LLM agents in such scenarios.
3.
We perform a holistic evaluation of LLM agents in Self-play and Cross-play settings, offering a detailed comparison with RL baselines and showcasing their potential as Coordination Agents.
4.
We investigate Environment Comprehension and Theory of Mind Reasoning as essential components of LLMs’ overall Joint Planning capabilities, highlighting their critical importance in coordination tasks.
2 LLM-Coordination Benchmark
2.1 Multi-turn Agentic Coordination
In the Multi-turn Agentic Coordination task, LLMs participate in end-to-end pure coordination games as agents, where the best strategy for all participating agents is to cooperate. LLMs under test are plugged into coordination frameworks with components like memory modules and grounding modules to act in complete games. These LLM agents can then be partnered with any policies or agents to complete the games.
Our LLM-Coordination benchmark includes 4 pure coordination games, Hanabi Challenge (Bard et al., 2020), Overcooked-AI (Carroll et al., 2019a), Collab Capture and Collab Escape.
Hanabi Challenge: In Hanabi (Bard et al., 2020), players aim to assemble five sequences of cards in ascending order (1 through 5). A unique aspect of the game is that the players can only view their partner’s cards, not their own. This requires players to work collaboratively, utilizing reveal tokens to provide hints about the cards in their partner’s hand. These hints can be about either the color or the rank of the cards. For instance, using a single hint token, a player can indicate all cards of a certain rank in their partner’s hand. Hanabi serves as an exemplary Pure Coordination game, necessitating player cooperation to achieve optimal outcomes. Success in Hanabi hinges on the ability to understand partners’ perspectives, navigate decisions based on incomplete information, and engage in implicit communication, making it an excellent testing ground for coordination among agents.
Overcooked-AI: In the Overcooked-AI environment (Carroll et al., 2019a), two agents—Alice (Blue) and Bob (Green)—collaborate to cook and deliver onion soups. This environment includes a variety of layouts, each with its own arrangement and quantity of onion dispensers, plate dispensers, cookers, delivery zones, and countertops. To prepare a dish, agents are required to insert three onions into a cooker, initiating a cooking process that lasts 20 time steps. Upon completion, the soup must be plated and delivered to complete the task. Each layout presents unique challenges, emphasizing the need for agents to comprehend their surroundings, locate necessary resources, and synchronize their actions with their teammate for effective collaboration.
Collab Capture: Collab Capture involves two agents trying to capture an adversary in a maze of interconnected rooms. The rooms are connected by doors, which can be controlled through access buttons that can be found in other rooms. The agents’ task is to capture the adversary in the least amount of time using effective strategies, including cornering the adversary and manipulating the doors to enable their partner or disable the adversary.
Collab Escape: Collab Escape involves two agents trying to escape an adversary in a maze of interconnected rooms. They need to fix two generators (similar to the game Dead-by-Daylight (Dea, 2016)) located in different rooms to open an exit portal. The adversary tries to catch the agents, and the win condition is any one agent escaping. This game requires strategies like luring the adversary away from the partner, sacrificing for the partner’s safety, and manipulating the movement of the adversary.
2.2 Single-turn Coordination QA
The agentic coordination task paints a holistic picture of the abilities of LLMs as agents. To dive deeper into the specific strengths and weaknesses of LLMs, we develop the CoordinationQA Suite. Inspired by the idea of Unit Testing for evaluating AI agents Knott et al. (2021), we manually sampled edge cases from all 4 pure coordination games mentioned in Section 2.1. All of these edge cases necessitate agents to actively understand their current state, think about their partner’s intentions, and come up with the best plans for coordination. We then create a set of three types of questions for each scenario in our CoordinationQA Suite.
Environment Comprehension (EC) questions require LLMs to make indirect inferences about some aspect of their environment.
Theory of Mind Reasoning (ToM) questions challenge the LLMs to predict the intentions of their partners and probe about the requirements of their partners.
Joint Planning (JP) questions provide agents with the state/observation and ask them to predict the best next action for effective coordination. This question is essentially the same question that LLMs need to repeatedly solve when they act as agents.
All the questions were manually developed and labeled. We filtered out questions and scenarios that showed any ambiguity, leaving only questions that had clear optimal solutions. We generated a total of N=66 scenarios (25 from Overcooked, 28 from Hanabi, and 13 from the two Collab Games) and created 3 questions per scenario, resulting in 198 unique questions. The right side of Figure 1 demonstrates the sampling process for the three types of questions with an example from the game Overcooked. The selected scenario shows the Blue agent about to place their third onion in the cooker, and the green agent needs to figure out what to do next.
3 Cognitive Architecture for Coordination
We develop a LLM Agent architecture based on Sumers et al. (2023), which we dub Cognitive Architecture for Coordination (CAC), for multi-agent coordination. Using CAC we can easily plug and play a LLM agent, and pair it with any partner (e.g., another CAC agent, a human player, or other AI agents). The architecture consists of three key elements: Memory, Reasoning, and Grounding.
Memory Module includes (1) Long-Term Memory for storing the Game Description including the game’s rules, conventions, objectives and action space, (2) Working memory which consists a textual description of the current observation, and (3) Episodic Memory which is a list of previous actions selected by the agent. In the example shown in Figure 2 the Long Term Memory includes a description of the game of Hanabi, The Working memory includes observations including the current stack, partner’s hand, beliefs and knowledge of both agents and information regarding the available tokens, and cards, and the Episodic Memory includes the previously selected discards, plays and hints.
Reasoning Module is where the Large Language Model (LLM) is plugged into the framework. It takes the textual description in the working memory as input and generates the next best action based on the context. For coordination games like Hanabi that require a more sophisticated Theory of Mind reasoning, we add an auxiliary Theory of Mind Reasoning LLM whose sole responsibility is to interpret the partner agent’s actions and requirements. This extra reasoning is added to the working memory before passing to the primary LLM. Additionally, we also utilize a Self-verification LLM which verifies the safety of the selected action. In the Hanabi example in Figure 2, the reasoning module interprets the provided clue of ”Revealing Rank 1 Cards” and decides to play one of the pointed cards on the empty stacks. The generated action is passed to the grounding module.
Grounding Module is responsible for interfacing the reasoning and memory modules’ textual decision-making spaces with the actual game mechanics. Its primary task is to take the selected action from the reasoning module and translate it into game-compatible action(s). The exact implementation of the grounding module depends on the game in question; for example, in Overcooked-AI, the grounding module needs to convert high-level actions like ”pick up onion from o0.” into sequences of lower-level actions. On the other hand, in games like Hanabi, it just needs to match actions like ”Reveal Bob’s Red Color Cards” to their lower-level representations. The grounding module is also responsible for the secondary task of filtering out infeasible actions based on the context of the game.
4 Experiments
4.1 Agentic Coordination
4.1.1 Setup
We perform two types of experiments in agentic coordination: Self-Play and Cross-Play. In self-play settings, the participating agents are of the same type. In Cross-Play experiments, we pair agents with unseen partners, and they need to adapt their behavior to the actions of these new partners.
Self-play Baselines: For Overcooked we use Proximal Policy Optimization (Schulman et al., 2017) and Population-Based Training (Jaderberg et al., 2017) as baselines for comparison. These baselines were established by Carroll et al. (2019a). The Hanabi challenge has been extensively studied and solved using MARL methods. We use Bayesian Action Decoder (Bard et al., 2020), Simplified Action Decoder (Hu & Foerster, 2021), and Off-Belief Learning (Hu et al., 2021a) as baselines.
Cross-play Baselines: For Overcooked, we use a Behavior Cloning model trained on human data Carroll et al. (2019a) and a Proximal Policy Optimization (PPO) agent trained with the Human Behavior Cloning agent Carroll et al. (2019a) as baselines for comparison. We also use human proxies based on behavior cloning as unseen partners. For Hanabi, we use the Simplified Action Decoder (SAD) as a baseline. We pair our agents with the Off-Belief Learning (Hu et al., 2021a), which was trained to generate grounded policies and adapt to unseen partner agents.
Metrics: We measure the total score achieved by agents in Overcooked, where each delivery provides 20 points to both agents. In the case of Hanabi, the metric is the total number of cards that have been correctly arranged by the players.
4.1.2 Results and Analysis
Overcooked Layouts | |||||
Method | CR | AA | Ring | FC | CC |
PPO (Schulman et al., 2017) | |||||
PBT (Jaderberg et al., 2017) | |||||
CACGPT-4-turbo | |||||
CACGPT-3.5-turbo | |||||
CACMixtral8x7B |
Class | Method | Score |
RL | BAD (Foerster et al., 2019) | |
SAD (Hu & Foerster, 2021) | ||
OBL (Hu et al., 2021a) | ||
CAC | GPT-4-turbo | |
GPT-3.5-turbo | ||
Mixtral-8x7b |
Method | Score |
LLM+Self verif.+ToM | |
LLM+Self Verif. | |
LLM |
LLM Agents outperform or match state-of-the-art RL methods in coordination games that depend more on understanding the environment.
We observed that LLM agents (w. GPT-4-turbo) outperform or match the overall performance of RL methods across all layouts of Overcooked-AI. Table 1 presents the numerical scores attained by different agents when paired with a partner agent of the same type. This implies that LLM agents outdo RL agents that have been trained together through Self-play without any game-specific training or fine-tuning. It is, however, important to note that LLM agents are significantly slower and larger than RL models and are not fit for real-time use yet (latency (seconds) of with Chain-of-thought and without for GPT-4-turbo). Furthermore, other models we tested, GPT-3.5-turbo and Mixtral8x7b, are faster but fall short of the RL baselines. We also see positive results on the CollabCapture and CollabEscape games with CAC agents (w. GPT-4), achieving a 100% success rate. However, other LLMs are unable to crack CollabEscape (see Appendix D).
LLM agents struggle at effective planning when advanced Theory of Mind reasoning is required. In Hanabi Challenge, LLM agents seem to struggle compared to RL methods (see Table 2). Among all LLMs, GPT-4-turbo performs reasonably well, while other LLMs can barely complete the Hanabi games. We attribute this failure to two factors. First, there is little room for errors in Hanabi. Any misplay or mis-clue leads to the loss of a life token. Second, Hanabi requires much more complex Theory of Mind Reasoning compared to the Overcoked-AI environment. Each action requires agents to actively consider their partner’s beliefs, intentions, and how they would react to implicit communication.
In contrast, Overcooked is fully observable, and its action space consists of actions like pick up an onion from onion_dispenser_0 and place onion in cooker_0. Under most scenarios and layouts, LLMs only need to consider the next best steps based on the state of the environment. For example, We conduct an ablation study of removing partner inventory and location, which reveals minimum impact on overall performance (1 less delivery in Cramped Room and Asymmetric Advantage layouts each in 100 timesteps), showing that the primary challenge for LLMs in games like Overcooked is the Environment Comprehension ability.
LLM agents benefit from auxiliary reasoning modules in imperfection information games with low room for errors. Without the support of auxiliary modules, LLM agents seem to bomb (lose all three lives) in every game (see Table 3). To rectify this, our CAC framework incorporates auxiliary LLM-powered modules, including an Explicit Theory of Mind Reasoning (ToM) module and an Answer Verification (AV) module. The answer verification module is a game-changer in its ability to stop LLM hallucinations about environmental facts from causing fatal mistakes, thus reducing the chance of mis-plays. The ToM reasoning LLM delegates the responsibility of interpreting partner clues and understanding partner needs to different LLMs, allowing the primary LLM to focus on synthesizing the available information to plan the next action.
Overcooked Layouts | |||||
Method | CR | AA | Ring | FC | CC |
BC (Carroll et al., 2019a) | | | | | | | | | | |
PPOBC (Schulman et al., 2017) | | | | | | | | | | |
CACGPT-4-turbo111For CAC, we run a single trial from either position due to cost constraints. In Table 1 we have observed that the performance of CAC agents does not vary by more than one delivery. | | | | | | | | | | |
Method | Self-Play | Cross-Play w/ OBL-1 | Cross-Play w/ OBL-4 |
SAD (Hu & Foerster, 2021) | |||
CACGPT-4-turbo |
LLM Agents are robust to unseen partners. We use Overcooked-AI and the Hanabi challenge as testbeds to evaluate the performance of LLM agents when paired with unseen agents. This task is popularly known as Zero Shot Coordination. For experiments in Overcooked-AI, we pair our LLM agents as well as baselines with proxy-human agents. These proxy human agents are behavior cloning agents trained using human data by Carroll et al. (2019b). As shown in Table 4, we discover that LLM agents outperform both Behavior Cloning as well as PPO agents trained with human data.
For experiments in Hanabi, we pair our agents with Off-Belief Learning (OBL) agents (Hu et al., 2021a). OBL is a MARL strategy that generates grounded clues and actions and is the state-of-the-art method for both self-play and cross-play in Hanabi. OBL agents provide observation-grounded clues and collaborate well with humans. Therefore, we use them as unseen partners in our experiments. Table 5 shows that CAC agents score an average of 15.00 points with the OBL-1 agent compared to their self-play scores of 13.66. This indicates no degradation in coordination abilities with a new partner. The baseline RL method, Simplified Action Decoder (SAD) Hu & Foerster (2021), fails critically when paired with unseen OBL agents, even though it excels at self-play (22.00 points) due to self-play training.
MARL agents trained with self-play struggle when paired with unseen partners in common payoff tasks, because they converge to arbitrary policies that only the two partners involved in the self-play training understand (Carroll et al., 2019a; Bard et al., 2020). Since LLM agents haven’t been explicitly trained to play these games, they base their outputs on the provided textual observation and commonsense knowledge learned from pre-training, and thus are much more robust to unseen partners.
4.2 Coordination QA
Setup.
We assess the performance of 6 Families of Large Language Models (LLMs) Jiang et al. (2023; 2024); Touvron et al. (2023); Chiang et al. (2023); OpenAI (2023) across three dimensions: Environment Comprehension (EC), Theory of Mind Reasoning (ToM), and Joint Planning (JP). For each category, LLMs respond to multiple-choice questions (MCQs), with their responses evaluated against ground-truth answers through fuzzy string matching. To account for the variability in LLM responses, we conduct three trials per model. We also report a Random baseline.
Comparative Results of LLMs in Environment Comprehension, ToM Reasoning, vs. Joint Planning. In Figure 3, we see that most LLMs achieve their best results on the Environment Comprehension question. The best performing LLM GPT-4-turbo gets more than 80% Environment Comprehension Questions correct. The overall performance across LLMs drops on the more challenging Theory of Mind reasoning questions, but GPT-4-turbo is still competent, reaching a 54% accuracy. The overall accuracy of LLMs on Joint Planning questions is still significantly weak, with even the best LLM scoring less than 40%, indicating a large room for improvement in LLMs’ ability to perform coordination reasoning. Another cause for concern is that open-source LLMs perform abysmally at Joint Planning, with some models performing worse than a random baseline.
Variables | ||
ToM | 0.813 | 0.389 |
EC | 0.895 | 0.506 |
Impact of Environment Comprehension and ToM Reasoning abilities on Joint Planning. Having defined Joint Planning as the capacity of an agent to select the appropriate subsequent action based on available information, we argue that proficiency in Environment Comprehension and Theory of Mind Reasoning is crucial for adept Joint Planning, and LLMs that do well at these two will do well at JP. Correlation analysis of ToM and EC with JP across the data from Figure 6 reveals that ToM has a moderate positive correlation to JP, whereas EC shows a strong positive correlation.
5 Related Work
Multi-agent Coordination
In Game Theory, Pure Coordination games are situations where the payoff is commonly shared between all agents. In such situations, cooperating is the best strategy. Various benchmarks and games have been used to evaluate Multi-Agent Coordination abilities over the years including Multiparticle Environment Lowe et al. (2017), Overcooked-AI Carroll et al. (2019a), and the Hanabi Challenge Bard et al. (2020). The foundational work by Carroll et al. (2019a) emphasized the significance of incorporating human data for effective human-ai collaboration. Subsequent research on the Overcooked-AI challenge has pivoted towards enabling self-play-trained agents to coordinate seamlessly with humans within this environment. These studies employ various techniques, including self-play with past agent checkpoints (Strouse et al., 2021), centralized population entropy objectives (Zhao et al., 2023), open-ended objectives using graph theory (Li et al., 2023a), policy ensembles with context-aware mechanisms (Lou et al., 2023), and the incorporation of human biases as linear hidden rewards (Yu et al., 2023), to enhance the training and diversity of AI agents in different scenarios. On the Hanabi Challenge much effort has been made to learn grounded policies Hu et al. (2021b; a) over arbitrary conventions. Embodied environments usually set up in household environments have also been recently used to study multi-agent coordination (Puig et al., 2021; Jain et al., 2020; 2019; Gan et al., 2021). The Overwhelming majority of approaches to coordination problems have focused on utilizing and enhancing Reinforcement Learning methods to solve the problems of multi-agent coordination. In this work, we argue that Large Language Models are an alternative approach to these coordination problems as they show emergent reasoning abilities, demonstrate theory-of-mind-like abilities, and do not converge to policies that cause arbitrary joint interactions.
Planning and Reasoning with Large Language Models
Large Language Models (LLMs) have demonstrated remarkable capabilities of reasoning in natural language (OpenAI, 2023; Ouyang et al., 2022; Chiang et al., 2023), achieving state-of-the-art performance across a spectrum of verbal reasoning tasks. It was then shown that LLMs could be augmented with components like memory, tools, perception, and grounding to create agents that could interact with an external environment (the web, simulators, games, etc.) These LLM agents have shown to be capable of solving long-horizon tasks, playing complex games (Wu et al., 2023; Wang et al., 2023) and interacting with simulated embodied environments (Liang et al., 2022; Song et al., 2022). Zhang et al. (2023b) developed a modular agent framework that was capable of cooperating with partner agents in embodied spatial rearrangement problems, demonstrating increased efficiency through coordination. Zhang et al. (2023a) develop a specialized architecture that enables LLMs to play in the Overcooked-AI environment. Li et al. (2023b) evaluate and show emergent collaborative abilities of LLMs in gamified simulations. In contrast to existing works, our work focuses on evaluating language agents in established pure coordination games where coordination is not an optional efficiency enhancer but rather a necessity.
6 Conclusion
In this study, we evaluated and analyzed the current large language models in terms of their ability to reason and act in pure coordination games. We introduced the LLM-Coordination benchmark with its two tasks: 1. Agentic Coordination and 2. CoordinationQA. These settings allowed us to conduct holistic comparative studies of LLMs as agents as well as dive deeper into the fine-grained aspects of LLMs as coordination reasoners. We juxtaposed LLM agents with existing Multi-agent Reinforcement Learning agents, discussing the conditions in which LLMs thrive and fail. Finally, we discussed the Theory of Mind Reasoning and Environment Comprehension as prerequisites for coordination and evaluated existing LLMs on these two components.
Appendix A Overcooked Implementation Details
A.1 Game and Layout Description
We use a general game description that explains the rules and objectives of overcooked. Since each layout has a different number of locations, like onion dispensers and cookers, we include a succinct description of each environment , which includes how many instances of particular facilities there are. For environments that include partitions, we mention which partition each of the agents is situated in and what facilities that agents can access. In addition, we also mentioned the shape of the environment.
A.2 State Description
The State is represented in natural language in the working memory, which can be processed by a Large Language Model (LLM). The state includes variables that fully represent the necessary details of the layout as well as the players. The information provided in is equivalent to what would be accessible to a Reinforcement Learning (RL) agent in the form of state representations. The following information is included in :
Objects Held by Each Player
The state description begins by detailing the inventories and of Alice and Bob, respectively. Each inventory (where ) can contain one of the following items: {”onion”, ”plate”, ”cooked soup”}. This inventory information is translated into natural language and incorporated into in the format: “I am holding . Bob is holding .” Such information is vital for inferring the likely subsequent actions of the partner agent.
Location of the Agent Controlled by LLM:
Given the limitations of Large Language Models (LLMs) in interpreting grid-based spatial information, we opt to provide processed location data to the LLM. For each agent (where ), and for each location of interest denoted as loc, we calculate the distance as the number of steps required to reach loc from using the shortest available path. The state description then includes this processed location information in the format: “loc is units away.” Here, loc can represent various points of interest such as onion dispensers, plate dispensers, cookers, delivery areas, kitchen counters, or shared counters. If a location is either inaccessible or blocked by another agent, this is explicitly stated in . For example, if a location is blocked by Bob, it would be stated as “loc is blocked by Bob.” To distinguish between the location information relevant to each agent, prefixes the respective sections with “Your location information:” for the agent controlled by the LLM and “Bob’s location information:” for the partner agent.
Cooker Information
The state description also incorporates information about the cooker, which is central to the gameplay strategy. Specifically, for each cooker , includes the number of onions currently in the pot. Additionally, provides the operational state of the cooker, denoted as , which can be either ”Off” or ”On”. Lastly, the current condition of the soup in the cooker is represented by , which can take one of the following values: ”Cooking”, ”Cooked”, or ”Not Started”. Thus, the information for cooker is formatted as: “ has onions. is . Soup in is .”
Kitchen Counter Information
The state description includes information about kitchen counters, which are primarily used for temporary object storage. Specifically, identifies the closest empty kitchen counter and the set of all counters currently holding an object.
Shared Counter Information
Shared counters serve as specialized kitchen counters for object transfer between agents. For each shared counter , includes the status for , as “ is empty” or “ contains onion,” to offer a complete environmental overview. Unlike kitchen counters, where only the closest empty counter is mentioned, all empty shared counters are mentioned.
Appendix B Hanabi Implementation Details
B.1 Game Description
We structure the game description of Hanabi into the overall objective, the rules of the game, and list of conventions based on the H-group conventions han (2024). We do not use advanced conventions like Chop Cards but stick to basic conventions about the card layout, play clues, and save clues.
B.2 State Description
The state description includes the current Stack , the player’s knowledge of their cards (updated based on clues), the partner agent’s cards , the partner agent’s knowledge of their cards (updated based on previous clues), each card in the discard pile , the remaining Life Tokens , and reveal tokens and the remaining Deck Size . We also precalculate the next card that goes on each stack since LLMs frequently fail to count which card should go next on each stack.
Appendix C Examples of prompts of LLMs used in CAC framework
C.1 Action Generator
The action generator generates a brief explanation and selected the next action to be played.
C.2 Theory of Mind Reasoning LLM
The Theory of Mind Reasoning LLM interprets partner actions and clarifies partner requirements.
C.3 Self Verification LLM
The prompt for the Self Verification LLM is provided as a system prompt. We observed that LLMs were more likely to act as strict verifiers when the prompt is provided as a system prompt.
Appendix D Results of Different LLMs on CollabCapture and CollabEscape
Table 7 summarizes the performance of the three LLMs GPT-4-turbo, GPT-3.5-turbo and Mixtral-8x7b on CollabGames - Collab Capture and Collab Escape. GPT-4-turbo achieves a 100% success rate in both games. In general, CollabEscape was a more difficult game to solve than CollabGames since agents needed to perform sacrificial moves and lure the killer away, requiring more advanced Theory of Mind inference.
Collab Capture | Collab Escape | |||
CAC model | Capture Rate | Turns to Capture | Escape Rate | Turns to Escape |
GPT-4-turbo | 1.00 | 3.99 | 1.00 | 24.67 |
GPT-3.5-turbo | 0.50 | 8.49 | 0.00 | N.A. |
Mixtral-8x7b | 0.75 | 3.88 | 0.00 | N.A. |
Greedy Baseline | 0.50 | 6.00 | 0.00 | N.A. |
Appendix E Generating Questions for CoordinationQA
E.1 Environment Comprehension Questions
The Environment Comprehension (EC) questions are indirect formulations regarding spatial aspects of the layout. In order for an agent to correctly answer an EC question, they must have an understanding of the dynamic details of the current state, the rules of the game, and exhibit spatial awareness. As such, when creating the EC questions, we carefully comb through a given scenario in search of salient points to probe an agent’s understanding of the given environment. Some examples include:
E.2 Theory of Mind Reasoning Questions
There are two primary question types in Hanabi for ToM Reasoning questions. In the first type, we ask the LLM about what information the partner agent needs, while in the second type, we ask it to make inferences about the partner agent’s last action. For all games apart from Hanabi, the ToM questions ask the models to predict the next intended action of the partner agent
E.2.1 Hanabi Question Type-1
E.2.2 Hanabi Question Type-2
E.2.3 Other Games
E.3 Joint Planning Questions
Joint planning questions are effectively the same questions that the LLM solves when they are part of an agentic framework. For each scenario, we ask the LLM to answer the question: ”What is the best next action?”.

### v3

@@
 Abstract
-The emergent reasoning and Theory of Mind (ToM) abilities demonstrated by Large Language Models (LLMs) make them promising candidates for developing coordination agents.
-In this study, we introduce a new LLM-Coordination Benchmark aimed at a detailed analysis of LLMs within the context of Pure Coordination Games, where participating agents need to cooperate for the most gain.
-This benchmark evaluates LLMs through two distinct tasks: (1) Agentic Coordination, where LLMs act as proactive participants for cooperation in 4 pure coordination games; (2) Coordination Question Answering (QA), where LLMs are prompted to answer 198 multiple-choice questions from the 4 games for evaluation of three key reasoning abilities: Environment Comprehension, ToM Reasoning, and Joint Planning.
-Furthermore, to enable LLMs for multi-agent coordination, we introduce a Cognitive Architecture for Coordination (CAC) framework that can easily integrate different LLMs as plug-and-play modules for pure coordination games.
-Our findings indicate that LLM agents equipped with GPT-4-turbo achieve comparable performance to state-of-the-art reinforcement learning methods in games that require commonsense actions based on the environment.
-Besides, zero-shot coordination experiments reveal that, unlike RL methods, LLM agents are robust to new unseen partners.
-However, results on Coordination QA show a large room for improvement in the Theory of Mind reasoning and joint planning abilities of LLMs.
-The analysis also sheds light on how the ability of LLMs to understand their environment and their partner’s beliefs and intentions plays a part in their ability to plan for coordination.
-Our code is available at https://github.com/eric-ai-lab/llm_coordination.
+Large Language Models (LLMs) have demonstrated emergent common-sense reasoning and Theory of Mind (ToM) capabilities, making them promising candidates for developing coordination agents.
+This study introduces the LLM-Coordination Benchmark, a novel benchmark for analyzing LLMs in the context of Pure Coordination Settings, where agents must cooperate to maximize gains.
+Our benchmark evaluates LLMs through two distinct tasks.
+The first is Agentic Coordination, where LLMs act as proactive participants in four pure coordination games.
+The second is Coordination Question Answering (CoordQA), which tests LLMs on 198 multiple-choice questions across these games to evaluate three key abilities: Environment Comprehension, ToM Reasoning, and Joint Planning.
+Results from Agentic Coordination experiments reveal that LLM-Agents excel in multi-agent coordination settings where decision-making primarily relies on environmental variables but face challenges in scenarios requiring active consideration of partners’ beliefs and intentions.
+The CoordQA experiments further highlight significant room for improvement in LLMs’ Theory of Mind reasoning and joint planning capabilities.
+Zero-Shot Coordination (ZSC) experiments in the Agentic Coordination setting demonstrate that LLM agents, unlike RL methods, exhibit robustness to unseen partners.
+These findings indicate the potential of LLMs as Agents in pure coordination setups and underscore areas for improvement.
+Code Available at https://github.com/eric-ai-lab/llm_coordination.
+LLM-Coordination: Evaluating and Analyzing Multi-agent Coordination Abilities in Large Language Models
 1 Introduction
@@
 These games require agents to reason about their environment and plan while considering the beliefs and intentions of their partners.
-Recently, Large Language Models (LLMs) have demonstrated emergent planning abilities in both physical and virtual settings (Raman et al., 2022; Wang et al., 2023; Wu et al., 2023), impressive reasoning capabilities (Wei et al., 2022), and the hints of a Theory of Mind (Kosinski, 2023) making them promising candidates for developing coordination agents.
+Recently, Large Language Models (LLMs) have demonstrated emergent planning abilities in both physical and virtual settings (Raman et al., 2022; Wang et al., 2023a; Wu et al., 2023), impressive reasoning capabilities (Wei et al., 2022), and the hints of a Theory of Mind (Kosinski, 2023) making them promising candidates for developing coordination agents.
 Previous works have explored the use of LLMs for developing collaborative agents, yet the requisite conditions, strengths, and limitations of LLMs in coordination games remain unclear.
 In this study, we intend to bridge the gap by performing a comprehensive evaluation and analysis of the multi-agent coordination abilities of LLMs.
-Therefore, we introduce a new LLM-Coordination Benchmark featuring two task settings for pure coordination games: 1.
+Therefore, we introduce the LLM-Coordination Benchmark featuring two task settings for pure coordination games: 1.
 Agentic Coordination and 2.
@@
 In Agentic Coordination, LLMs are scaffolded with components that allow them to act within actual game environments, providing a holistic evaluation of the competencies of LLMs to act as coordination agents.
-In CoordinationQA, LLMs have to answer a curated set of questions about edge-case scenarios drawn from coordination games where agents are required actively cooperate with their partners.
+In CoordinationQA, LLMs have to answer a curated set of questions about edge-case scenarios drawn from coordination games where agents need to actively cooperate with their partners.
 The benchmark includes four collaborative games, providing a comprehensive analysis platform.
-To enable LLMs for multi-agent coordination, we present a Cognitive Architecture for Coordination (CAC) framework that facilitates LLM interaction with game environments in a plug-and-play approach.
-CAC translates game elements into textual formats and leverages auxiliary LLMs for improved coordination to enable effective multi-agent collaboration.
-Our Experiments on the Agentic Coordination task with CAC reveal that Large Language Models are capable of understanding the game objectives, generating coherent reasoning for their next actions, and coordinating with partners across all coordination games.
-They exhibit these competencies without any training, fine-tuning, or few-shot examples.
-A comparative analysis on the Agentic Coordination task reveals that LLM agents outperform Multi-agent RL solutions in games that require minimum theory-of-mind reasoning and focus on taking commonsense actions based on the environment.
-However, they struggle in more complex settings where agents need to actively consider their partner’s beliefs and intentions.
+Unlike studies on multi-LLM frameworks Hong et al.
+(2023); Qian et al.
+(2024); Li et al.
+(2023a), which focus on orchestrating multiple LLMs to solve tasks, our benchmark assesses the innate ability of individual LLMs to understand and act within pure coordination scenarios where cooperation is essential.
+Our experiments in the Agentic Coordination setting reveal that Large Language Models are competent at understanding the game objectives, generating coherent reasoning for their next actions, and coordinating with partners across all coordination games.
+They exhibit these behaviors without any training, fine-tuning, or few-shot examples.
+A comparative analysis reveals that LLM agents match or outperform RL baselines in games where optimal decision-making can be done by observing environment variables and positions (e.g., Overcooked).
+However, they struggle in settings where agents need to actively consider their partner’s beliefs and intentions (e.g., Hanabi).
 We also observe that LLM agents are capable of collaborating with new partners, unlike self-play MARL methods (Carroll et al., 2019a; Bard et al., 2020) that fail to adapt to unseen agents.
@@
 This suite is designed to dissect the capabilities of LLMs in single-turn reasoning within coordination games, focusing on three key areas: Joint Planning, Theory of Mind (ToM), and Environment Comprehension.
-Joint Planning evaluates LLMs’ decision-making for optimal long-term outcomes, ToM questions probe the understanding of partner agents’ intentions and needs, and Environment Comprehension assesses knowledge of game rules and dynamics.
-Our findings on CoordinationQA show a marked performance gap between GPT-4-turbo and other LLMs across three question types.
-LLMs are most proficient in Environment Comprehension, indicating they understand game rules and states well.
+Joint Planning (JP) evaluates LLMs’ planning abilities for optimal coordination, ToM questions probe their understanding of partner agents’ intentions and needs, and Environment Comprehension (EC) assesses their inference of environment details, rules and objectives.
+First, Our findings on CoordinationQA show a marked performance gap between GPT-4-turbo and other LLMs across three question types.
+Secondly, LLMs are most proficient in Environment Comprehension, indicating they understand the rules and environment states well.
 However, they face significant challenges in Theory of Mind Reasoning, with difficulty inferring others’ intentions and needs.
@@
 These results highlight LLMs’ limited reliability and effectiveness as coordination partners.
-Additionally, our analysis shows a moderate to strong correlation between Theory of Mind Reasoning, Environment Comprehension, and Joint Planning performance, underscoring their role in effective coordination.
-In summary, our contributions are four-fold:
+Correlation analysis between LLMs’ performance on CoordinationQA and their performance on agentic coordination setting further highlights their strengths in environmental reasoning but exposes significant weaknesses in Theory of Mind inference and Joint Planning capabilities.
+In summary, our contributions are threefold:
 1.
@@
 2.
-We develop the plug-and-play Cognitive Architecture for Coordination Framework to enable LLMs to effectively participate in complex, partially observable coordination games like Hanabi, demonstrating the first zero-shot application for LLM agents in such scenarios.
+We perform a holistic evaluation of LLM agents in Self-play and Cross-play settings, offering a detailed comparison with RL baselines and showcasing their potential as Coordination Agents.
 3.
-We perform a holistic evaluation of LLM agents in Self-play and Cross-play settings, offering a detailed comparison with RL baselines and showcasing their potential as Coordination Agents.
-4.
-We investigate Environment Comprehension and Theory of Mind Reasoning as essential components of LLMs’ overall Joint Planning capabilities, highlighting their critical importance in coordination tasks.
+We investigate Environment Comprehension, Theory of Mind Reasoning, and Joint Planning as essential components of LLMs’ overall coordination capabilities, highlighting their critical importance in pure coordination setups.
+2 Related Work
+Multi-agent Coordination.
+Pure Coordination games in game theory are scenarios where agents share the payoffs, and cooperation is the optimal strategy.
+Benchmarks like the Multiparticle Environment Lowe et al.
+(2017), Overcooked-AI Carroll et al.
+(2019a), and the Hanabi Challenge Bard et al.
+(2020) evaluate multi-agent coordination.
+Carroll et al.
+(2019a) highlighted the importance of human data for effective Human-AI collaboration.
+Subsequent Overcooked-AI research focuses on aligning self-play-trained agents with humans using techniques such as self-play with past checkpoints (Strouse et al., 2021), population entropy objectives (Zhao et al., 2023), graph-theoretic objectives (Li et al., 2023c), policy ensembles (Lou et al., 2023), and integrating human biases (Yu et al., 2023).
+In the Hanabi Challenge, efforts aim to learn grounded policies over arbitrary conventions Hu et al.
+(2021b, a).
+While most solutions enhance RL methods for coordination, we propose that LLMs offer an alternative due to their emergent reasoning and theory-of-mind-like abilities, avoiding arbitrary joint interactions.
+Planning and Reasoning with Large Language Models.
+LLMs have shown remarkable natural language reasoning abilities (OpenAI, 2023; Ouyang et al., 2022; Chiang et al., 2023), achieving state-of-the-art results in verbal reasoning tasks.
+Augmented with components like memory and tools, LLMs can interact with external environments, solving long-horizon tasks and playing complex games (Wu et al., 2023; Wang et al., 2023a; Liang et al., 2022; Song et al., 2022).
+Guided by Cognitive Architectures for Language Agents (Sumers et al., 2023) as a design principle for agent design, we experiment with advanced reasoning strategies such as ReAct Yao et al.
+(2023), Self-Verification Weng et al.
+(2023), and Self-Consistency Wang et al.
+(2023b) to enhance LLM reasoning.
+These strategies establish strong baseline performance for our Language Agent implementations.
+Multi-agent LLMs.
+Recent studies have explored LLMs in multi-agent cooperation settings.
+Zhang et al.
+(2023b) developed a modular agent framework for spatial rearrangement tasks.
+Zhang et al.
+(2023a) introduced an architecture enabling LLMs to play Overcooked-AI.
+Shi et al.
+(2023) demonstrated positive zero-shot coordination in Avalon using code-driven reasoning.
+Li et al.
+(2023d) showed emergent collaborative abilities of LLMs in simulations, while Li et al.
+(2023b) investigated theory-of-mind inference using explicit belief representations.
+Xu et al.
+(2024) perform an analysis of LLMs in communication games Xu et al.
+(2023) analyzed LLM cognitive abilities through games, highlighting benefits of probabilistic modeling.
+In contrast, our research rigorously evaluates LLM agents’ coordination abilities in established pure coordination games, where coordination is essential.
+Our setting only includes scenarios where agents must fully cooperate with each other with no competitive incentives.
+We also conduct a fine-grained, component-level analysis to understand the intricacies of LLMs’ coordination capabilities.
 3 LLM-Coordination Benchmark
@@
 In the Multi-turn Agentic Coordination task, LLMs participate in end-to-end pure coordination games as agents, where the best strategy for all participating agents is to cooperate.
-LLMs under test are plugged into coordination frameworks with components like memory modules and grounding modules to act in complete games.
+We only consider pure coordination scenarios with no competitive incentives.
+LLMs under test are plugged into coordination frameworks with memory and the ability to act in complete games.
 These LLM agents can then be partnered with any policies or agents to complete the games.
-Our LLM-Coordination benchmark includes 4 pure coordination games, Hanabi Challenge (Bard et al., 2020), Overcooked-AI (Carroll et al., 2019a), Collab Capture and Collab Escape.
-Hanabi Challenge: In Hanabi (Bard et al., 2020), players aim to assemble five sequences of cards in ascending order (1 through 5).
+Our LLM-Coordination benchmark includes 4 pure coordination games: Hanabi Challenge (Bard et al., 2020), Overcooked-AI (Carroll et al., 2019a), and Collab Capture and Collab Escape (inspired by the Pursuit-Evasion problem).
+These games were carefully selected for their ability to isolate and highlight specific coordination challenges, providing controlled environments that allow the analysis of pure coordination scenarios without too much emphasis on other reasoning challenges.
+While LLMs are versatile and capable of addressing a wide range of tasks, these well-studied settings offer established benchmarks, clear metrics, and reproducible scenarios that are particularly suited for examining coordination abilities of participating agents.
+Hanabi Challenge.
+In Hanabi (Bard et al., 2020), players aim to assemble five sequences of cards in ascending order (1 through 5), each sequence dedicated to a different color: purple, red, blue, yellow, and green.
 A unique aspect of the game is that the players can only view their partner’s cards, not their own.
@@
 These hints can be about either the color or the rank of the cards.
-For instance, using a single hint token, a player can indicate all cards of a certain rank in their partner’s hand.
+For instance, using a single reveal token, a player can indicate all cards of a certain rank in their partner’s hand.
+Once a player has an idea about which card they have, they can choose to play the card on the stack.
+If the card is correct, they get a point.
+Otherwise, players lose a collective life token.
+The loss of all 3 life tokens leads to the end of the game.
 Hanabi serves as an exemplary Pure Coordination game, necessitating player cooperation to achieve optimal outcomes.
 Success in Hanabi hinges on the ability to understand partners’ perspectives, navigate decisions based on incomplete information, and engage in implicit communication, making it an excellent testing ground for coordination among agents.
-Overcooked-AI: In the Overcooked-AI environment (Carroll et al., 2019a), two agents—Alice (Blue) and Bob (Green)—collaborate to cook and deliver onion soups.
+Overcooked-AI.
+In the Overcooked-AI environment (Carroll et al., 2019a), two agents—Alice (Blue) and Bob (Green)—collaborate to cook and deliver onion soups.
 This environment includes a variety of layouts, each with its own arrangement and quantity of onion dispensers, plate dispensers, cookers, delivery zones, and countertops.
@@
 Each layout presents unique challenges, emphasizing the need for agents to comprehend their surroundings, locate necessary resources, and synchronize their actions with their teammate for effective collaboration.
-Collab Capture: Collab Capture involves two agents trying to capture an adversary in a maze of interconnected rooms.
+Collab Capture.
+Collab Capture involves two agents trying to capture an adversary in a maze of interconnected rooms.
 The rooms are connected by doors, which can be controlled through access buttons that can be found in other rooms.
-The agents’ task is to capture the adversary in the least amount of time using effective strategies, including cornering the adversary and manipulating the doors to enable their partner or disable the adversary.
-Collab Escape: Collab Escape involves two agents trying to escape an adversary in a maze of interconnected rooms.
+The agents’ task is to capture the adversary in the least amount of time using effective strategies.
+To evaluate different coordination strategies between agents, we design four scenarios by controlling the states of the doors (open or closed) within the layout shown in figure 5.
+These scenarios highlight various coordination challenges, such as trapping the adversary through precise positioning, enabling a teammate by prioritizing door control over direct pursuit, and strategically restricting the adversary’s movement to facilitate capture.
+(see Appendix E.1 for more details.)
+Collab Escape.
+Collab Escape involves two agents trying to escape an adversary in a maze of interconnected rooms.
 They need to fix two generators (similar to the game Dead-by-Daylight (Dea, 2016)) located in different rooms to open an exit portal.
 The adversary tries to catch the agents, and the win condition is any one agent escaping.
-This game requires strategies like luring the adversary away from the partner, sacrificing for the partner’s safety, and manipulating the movement of the adversary.
+To evaluate coordination strategies in Collab Escape, we develop two scenarios by varying the initial proximity of agents to the adversary and generators in the layout shown in figure 6.
+Depending on their proximity to the adversary/generators, players need to apply strategies such as luring the adversary away from the partner, choosing to continue fixing the generators while sacrificing for the partner’s safety and manipulating the movement of the adversary (see Appendix E.2 for more details.)
 3.2 Single-turn Coordination QA
@@
 We then create a set of three types of questions for each scenario in our CoordinationQA Suite.
-Environment Comprehension (EC) questions require LLMs to make indirect inferences about some aspect of their environment.
-Theory of Mind Reasoning (ToM) questions challenge the LLMs to predict the intentions of their partners and probe about the requirements of their partners.
+Environment Comprehension (EC) questions require LLMs to make indirect inferences about some aspect of their environment (See Appendix D.1).
+The questions cover details of the layouts, implications of current observations, and counts of artifacts.
+Theory of Mind Reasoning (ToM) questions challenge the LLMs to predict the intentions of their partners and probe about the requirements of their partners (See Appendix D.2).
 Joint Planning (JP) questions provide agents with the state/observation and ask them to predict the best next action for effective coordination.
-This question is essentially the same question that LLMs need to repeatedly solve when they act as agents.
+This question is essentially the same question that LLMs need to repeatedly solve when they act as agents (See Appendix D.3).
 All the questions were manually developed and labeled.
-We filtered out questions and scenarios that showed any ambiguity, leaving only questions that had clear optimal solutions.
+We filtered out questions and scenarios that showed any ambiguity, leaving only questions that had clear, optimal solutions.
 We generated a total of N=66 scenarios (25 from Overcooked, 28 from Hanabi, and 13 from the two Collab Games) and created 3 questions per scenario, resulting in 198 unique questions.
@@
 The selected scenario shows the Blue agent about to place their third onion in the cooker, and the green agent needs to figure out what to do next.
-3 Cognitive Architecture for Coordination
-We develop a LLM Agent architecture based on Sumers et al.
-(2023), which we dub Cognitive Architecture for Coordination (CAC), for multi-agent coordination.
-Using CAC we can easily plug and play a LLM agent, and pair it with any partner (e.g., another CAC agent, a human player, or other AI agents).
-The architecture consists of three key elements: Memory, Reasoning, and Grounding.
-Memory Module includes (1) Long-Term Memory for storing the Game Description including the game’s rules, conventions, objectives and action space, (2) Working memory which consists a textual description of the current observation, and (3) Episodic Memory which is a list of previous actions selected by the agent.
-In the example shown in Figure 2 the Long Term Memory includes a description of the game of Hanabi, The Working memory includes observations including the current stack, partner’s hand, beliefs and knowledge of both agents and information regarding the available tokens, and cards, and the Episodic Memory includes the previously selected discards, plays and hints.
-Reasoning Module is where the Large Language Model (LLM) is plugged into the framework.
-It takes the textual description in the working memory as input and generates the next best action based on the context.
-For coordination games like Hanabi that require a more sophisticated Theory of Mind reasoning, we add an auxiliary Theory of Mind Reasoning LLM whose sole responsibility is to interpret the partner agent’s actions and requirements.
-This extra reasoning is added to the working memory before passing to the primary LLM.
-Additionally, we also utilize a Self-verification LLM which verifies the safety of the selected action.
-In the Hanabi example in Figure 2, the reasoning module interprets the provided clue of ”Revealing Rank 1 Cards” and decides to play one of the pointed cards on the empty stacks.
-The generated action is passed to the grounding module.
-Grounding Module is responsible for interfacing the reasoning and memory modules’ textual decision-making spaces with the actual game mechanics.
-Its primary task is to take the selected action from the reasoning module and translate it into game-compatible action(s).
-The exact implementation of the grounding module depends on the game in question; for example, in Overcooked-AI, the grounding module needs to convert high-level actions like ”pick up onion from o0.” into sequences of lower-level actions.
-On the other hand, in games like Hanabi, it just needs to match actions like ”Reveal Bob’s Red Color Cards” to their lower-level representations.
-The grounding module is also responsible for the secondary task of filtering out infeasible actions based on the context of the game.
-4 Experiments
+See Appendix D for examples of questions and the templates used to formulate these questions.
+4 Experimental Setup
 4.1 Agentic Coordination
-4.1.1 Setup
 We perform two types of experiments in agentic coordination: Self-Play and Cross-Play.
@@
 In Cross-Play experiments, we pair agents with unseen partners, and they need to adapt their behavior to the actions of these new partners.
-Self-play Baselines: For Overcooked we use Proximal Policy Optimization (Schulman et al., 2017) and Population-Based Training (Jaderberg et al., 2017) as baselines for comparison.
+4.1.1 LLM Agents
+To allow LLMs to play multi-turn games, we scaffold them with an agentic framework based on Cognitive Architectures for Language Agents Sumers et al.
+(2023).
+The framework includes three parts: Memory, Reasoning, and Grounding.
+Memory includes (1) Long-Term Memory for storing the Game Description, including the game’s rules, conventions, objectives, and action space, (2) Working memory, which consists of a textual description of the current observation, and (3) Episodic Memory which is a list of previous actions selected by the agent.
+Reasoning is where the Large Language Model (LLM) is plugged into the framework.
+It takes the textual description from the Memory as input and generates the next action based on the context.
+The LLM reasons about the current state and then selects an action from the list of available actions in natural language.
+Self-Verification: For the coordination game Hanabi, there is a low margin for error as any misplays lead to the loss of life tokens, and the loss of all three life tokens subsequently results at the end of the game.
+We thus supplement the reasoning process in Hanabi with Answer-Verification Weng et al.
+(2023), where the LLM is re-prompted to confirm that the action it generated is appropriate and does not lead to fatal errors.
+ToM-Reasoning: We also demonstrate the positive impact of a Theory of Mind Reasoning step prior to generating the next action for Hanabi and CollabEscape, which benefit from this intermediate step.
+In the ToM reasoning step, the LLM generates an interpretation of their partner’s actions or current position before generating the next action to explicitly capture the belief inference process.
+We do not test with additional ToM reasoning on Overcooked due to significant latency and cost constraints, with marginal benefits.
+Finally, the Grounding process translates the natural language action generated by the reasoning module into game-compatible action(s).
+The exact implementation of the grounding module depends on the game in question; for example, in Overcooked-AI, the grounding module needs to convert high-level actions like "pick up onion from o0." into sequences of lower-level actions.
+On the other hand, in games like Hanabi, the Grounding needs to match actions like "Reveal Bob’s Red Color Cards" to their lower-level representations.
+The Grounding process is also responsible for filtering out infeasible actions based on the context of the game (See Appendices A, B for more details.)
+There are no prompt or setup differences for LLM Agents based on Cross-play or Self-play.
+We use the LLMs gpt-4-0125-preview, GPT-3.5-turbo-0125, Mixtral 8x7B, and GPT-4o for agentic evaluation studies.
+| Agent | CR | AA | Ring | FC | CC |
+| Self-Play (PPO) | |||||
+| PBT | |||||
+| GPT-3.5-turbo | |||||
+| Mixtral8x7B | |||||
+| GPT-4o | |||||
+| GPT-4-turbo |
+| Collab Escape | Collab Capture | |||
+| Agent | Capture Rate | Avg. Turns | Escape Rate | Avg. Turns |
+| GPT-4-turbo | 0.83 | 4.60 | 1.00 | 3.5 |
+| – (w/out ToM Reasoning) | 0.50 | 2.00 | 1.00 | 4.75 |
+| GPT-4o | 0.67 | 4.00 | 1.00 | 7.17 |
+| GPT-3.5-turbo | 0.33 | 2.50 | 0.67 | 8.38 |
+| Mixtral-8x7b | 0.50 | 7.67 | 0.92 | 7.55 |
+| Greedy Baseline | 0.00 | N.A. | 0.50 | 6.00 |
+4.1.2 MARL Agents
+Self-play MARL Baselines: For Overcooked we use Proximal Policy Optimization (Schulman et al., 2017) and Population-Based Training (Jaderberg et al., 2017) as baselines for comparison.
 These baselines were established by Carroll et al.
 (2019a).
-The Hanabi challenge has been extensively studied and solved using MARL methods.
-We use Bayesian Action Decoder (Bard et al., 2020), Simplified Action Decoder (Hu & Foerster, 2021), and Off-Belief Learning (Hu et al., 2021a) as baselines.
-Cross-play Baselines: For Overcooked, we use a Behavior Cloning model trained on human data Carroll et al.
+For the Hanabi challenge, we use Bayesian Action Decoder (BAD) (Bard et al., 2020), Simplified Action Decoder (SAD) (Hu and Foerster, 2021), and Off-Belief Learning (Hu et al., 2021a) as MARL baselines for Hanabi.
+All three baselines achieve near-perfect performance in Self-play.
+Cross-play MARL Baselines: For Overcooked, we use a Behavior Cloning model trained on human data Carroll et al.
 (2019a) and a Proximal Policy Optimization (PPO) agent trained with the Human Behavior Cloning agent Carroll et al.
 (2019a) as baselines for comparison.
-We also use human proxies based on behavior cloning as unseen partners.
-For Hanabi, we use the Simplified Action Decoder (SAD) as a baseline.
-We pair our agents with the Off-Belief Learning (Hu et al., 2021a), which was trained to generate grounded policies and adapt to unseen partner agents.
-Metrics: We measure the total score achieved by agents in Overcooked, where each delivery provides 20 points to both agents.
+We also report Hidden-Utility Self-play (HSP) Yu et al.
+(2023) as a baseline.
+We use human proxies based on behavior cloning as unseen partners.
+For Hanabi, we use the Simplified Action Decoder (SAD), which is trained through self-play as a baseline.
+We pair our agents with Off-Belief Learning (Hu et al., 2021a), which was trained to generate grounded policies and adapt to unseen partner agents.
+Metrics.
+We measure the total score achieved by agents in Overcooked, where each delivery provides 20 points to both agents.
 In the case of Hanabi, the metric is the total number of cards that have been correctly arranged by the players.
-4.1.2 Results and Analysis
-Overcooked Layouts | |||||
-Method | CR | AA | Ring | FC | CC |
-PPO (Schulman et al., 2017) | |||||
-PBT (Jaderberg et al., 2017) | |||||
-CACGPT-4-turbo | |||||
-CACGPT-3.5-turbo | |||||
-CACMixtral8x7B |
-Class | Method | Score |
-RL | BAD (Foerster et al., 2019) | |
-SAD (Hu & Foerster, 2021) | ||
-OBL (Hu et al., 2021a) | ||
-CAC | GPT-4-turbo | |
-GPT-3.5-turbo | ||
-Mixtral-8x7b |
-Method | Score |
-LLM+Self verif.+ToM | |
-LLM+Self Verif. | |
-LLM |
-LLM Agents outperform or match state-of-the-art RL methods in coordination games that depend more on understanding the environment.
-We observed that LLM agents (w.
-GPT-4-turbo) outperform or match the overall performance of RL methods across all layouts of Overcooked-AI.
-Table 1 presents the numerical scores attained by different agents when paired with a partner agent of the same type.
-This implies that LLM agents outdo RL agents that have been trained together through Self-play without any game-specific training or fine-tuning.
-It is, however, important to note that LLM agents are significantly slower and larger than RL models and are not fit for real-time use yet (latency (seconds) of with Chain-of-thought and without for GPT-4-turbo).
-Furthermore, other models we tested, GPT-3.5-turbo and Mixtral8x7b, are faster but fall short of the RL baselines.
-We also see positive results on the CollabCapture and CollabEscape games with CAC agents (w.
-GPT-4), achieving a 100% success rate.
-However, other LLMs are unable to crack CollabEscape (see Appendix D).
-LLM agents struggle at effective planning when advanced Theory of Mind reasoning is required.
-In Hanabi Challenge, LLM agents seem to struggle compared to RL methods (see Table 2).
-Among all LLMs, GPT-4-turbo performs reasonably well, while other LLMs can barely complete the Hanabi games.
-We attribute this failure to two factors.
-First, there is little room for errors in Hanabi.
-Any misplay or mis-clue leads to the loss of a life token.
-Second, Hanabi requires much more complex Theory of Mind Reasoning compared to the Overcoked-AI environment.
-Each action requires agents to actively consider their partner’s beliefs, intentions, and how they would react to implicit communication.
-In contrast, Overcooked is fully observable, and its action space consists of actions like pick up an onion from onion_dispenser_0 and place onion in cooker_0.
-Under most scenarios and layouts, LLMs only need to consider the next best steps based on the state of the environment.
-For example, We conduct an ablation study of removing partner inventory and location, which reveals minimum impact on overall performance (1 less delivery in Cramped Room and Asymmetric Advantage layouts each in 100 timesteps), showing that the primary challenge for LLMs in games like Overcooked is the Environment Comprehension ability.
-LLM agents benefit from auxiliary reasoning modules in imperfection information games with low room for errors.
-Without the support of auxiliary modules, LLM agents seem to bomb (lose all three lives) in every game (see Table 3).
-To rectify this, our CAC framework incorporates auxiliary LLM-powered modules, including an Explicit Theory of Mind Reasoning (ToM) module and an Answer Verification (AV) module.
-The answer verification module is a game-changer in its ability to stop LLM hallucinations about environmental facts from causing fatal mistakes, thus reducing the chance of mis-plays.
-The ToM reasoning LLM delegates the responsibility of interpreting partner clues and understanding partner needs to different LLMs, allowing the primary LLM to focus on synthesizing the available information to plan the next action.
-Overcooked Layouts | |||||
-Method | CR | AA | Ring | FC | CC |
-BC (Carroll et al., 2019a) | | | | | | | | | | |
-PPOBC (Schulman et al., 2017) | | | | | | | | | | |
-CACGPT-4-turbo111For CAC, we run a single trial from either position due to cost constraints.
-In Table 1 we have observed that the performance of CAC agents does not vary by more than one delivery.
-| | | | | | | | | | |
-Method | Self-Play | Cross-Play w/ OBL-1 | Cross-Play w/ OBL-4 |
-SAD (Hu & Foerster, 2021) | |||
-CACGPT-4-turbo |
-LLM Agents are robust to unseen partners.
-We use Overcooked-AI and the Hanabi challenge as testbeds to evaluate the performance of LLM agents when paired with unseen agents.
-This task is popularly known as Zero Shot Coordination.
-For experiments in Overcooked-AI, we pair our LLM agents as well as baselines with proxy-human agents.
-These proxy human agents are behavior cloning agents trained using human data by Carroll et al.
-(2019b).
-As shown in Table 4, we discover that LLM agents outperform both Behavior Cloning as well as PPO agents trained with human data.
-For experiments in Hanabi, we pair our agents with Off-Belief Learning (OBL) agents (Hu et al., 2021a).
-OBL is a MARL strategy that generates grounded clues and actions and is the state-of-the-art method for both self-play and cross-play in Hanabi.
-OBL agents provide observation-grounded clues and collaborate well with humans.
-Therefore, we use them as unseen partners in our experiments.
-Table 5 shows that CAC agents score an average of 15.00 points with the OBL-1 agent compared to their self-play scores of 13.66.
-This indicates no degradation in coordination abilities with a new partner.
-The baseline RL method, Simplified Action Decoder (SAD) Hu & Foerster (2021), fails critically when paired with unseen OBL agents, even though it excels at self-play (22.00 points) due to self-play training.
-MARL agents trained with self-play struggle when paired with unseen partners in common payoff tasks, because they converge to arbitrary policies that only the two partners involved in the self-play training understand (Carroll et al., 2019a; Bard et al., 2020).
-Since LLM agents haven’t been explicitly trained to play these games, they base their outputs on the provided textual observation and commonsense knowledge learned from pre-training, and thus are much more robust to unseen partners.
+For CollabEscape and CollabCapture, we report the success rate of escape or capture across multiple trials and the average turns to capture or escape.
 4.2 CoordinationQA
-Setup.
-We assess the performance of 6 Families of Large Language Models (LLMs) Jiang et al.
-(2023; 2024); Touvron et al.
+We assess the performance of 5 Families of Large Language Models (LLMs) Jiang et al.
+(2023, 2024); Touvron et al.
 (2023); Chiang et al.
@@
 To account for the variability in LLM responses, we conduct three trials per model.
+All models being tested are shown the same prompts.
 We also report a Random baseline.
-Comparative Results of LLMs in Environment Comprehension, ToM Reasoning, vs.
-Joint Planning.
+5 Discussion
+Zero-shot LLM Agents match or surpass trained RL methods in Environment-focused Coordination Problems.
+We observed that LLM agents (w.
+GPT-4-turbo) outperform or match the overall performance of RL methods across all layouts of Overcooked-AI.
+Table 1 presents the numerical scores attained by different agents when paired with a partner agent of the same type.
+This implies that LLM agents match RL agents that have been explicitly trained through Self-play without any game-specific training or fine-tuning.
+However, it is important to note that LLM agents are significantly slower and larger than RL models, making them unsuitable for real-time use at present.
+We also see positive results on the CollabCapture and CollabEscape games, with most LLMs being able to complete both challenges (see Table 2).
+| Agent | Score |
+| Bayesian Action Decoder | |
+| Simplified Action Decoder | |
+| Off-Belief Learning | |
+| GPT-4-turbo | |
+| - (w.o ToM Reasoning) | |
+| - (w.o ToM Reasoning & Verif.) | |
+| GPT-4o | |
+| GPT-3.5-turbo | |
+| Mixtral-8x7b |
+LLM agents struggle at effective planning when advanced Theory of Mind reasoning is required.
+In Hanabi Challenge, LLM agents seem to struggle compared to RL methods (see Table 3).
+GPT-4-turbo performs reasonably well, while other LLMs can barely complete the games.
+We attribute this failure to two factors.
+First, there is little room for errors in Hanabi.
+Any misplay leads to the loss of a life token.
+Second, Hanabi requires more complex Theory of Mind Reasoning compared to the Overcooked-AI environment.
+Each action requires agents to actively consider their partner’s beliefs, intentions, and how they would react to implicit communication.
+In contrast, Overcooked is fully observable, and its action space consists of actions like pick up an onion from onion_dispenser_0 and place onion in cooker_0.
+Under most scenarios and layouts, LLMs only need to consider the next best steps based on the state of the environment.
+We use correlation study to provide additional validation to these findings.
+We calculate the Pearson Correlation Coefficient () of the performance of the four LLMs (GPT-4-turbo, GPT-4o, GPT-3.5-turbo, and Mixtral 8x7b) on Agentic Coordination setup (Average score per game) vs.
+the score on CoordinationQA task.
+Figure 2 shows a high correlation between Environment Comprehension capabilities and Success at Overcooked, but a more moderate correlation between Theory of Mind Reasoning capabilities and Success at Overcooked.
+Conversely, in Hanabi, high success is strongly correlated with both Environment Comprehension and Theory of Mind Reasoning Abilities.
+In CollabEscape we see a higher correlation with ToM reasoning abilities compared to CollabCapture.
+This correlation study also connects and establishes a positive alignment between LLM-agent performance on the multi-turn agentic task and the single-turn CoordinationQA.
+| Method | CR | AA | Ring | FC | CC |
+| BC | | | | | | | | | | |
+| PPOBC | | | | | | | | | | |
+| HSP | - | | | | | - | | |
+| GPT-4-turbo111For GPT-4-turbo, we run a single trial from either position due to cost and time constraints. | | | | | | | | | | |
+| Method | Self-Play | Cross-Play w/ OBL-1 | Cross-Play w/ OBL-4 |
+| SAD | |||
+| GPT-4-turbo |
+Auxiliary reasoning strategies like Verification and ToM reasoning help LLMs reason for coordination.
+Adding an Answer Verification step significantly reduces fatal mistakes (wrong card plays) caused by LLM hallucinations.
+Without the support of the Verification step, LLM agents bomb (lose all three lives) before the end of the game in every trial.
+The ToM reasoning step separates the tasks of interpreting partner clues and generating actions, allowing the LLM to better synthesize available information for action planning.
+Table 3 shows the impact of ablating the verification and ToM reasoning steps from the LLM Agent.
+The ToM reasoning step is also useful in the CollabEscape (see table 2) game, as players need to actively consider what their partner needs and act sacrificially if needed.
+In CollabCapture, it shows a relatively low benefit since agents can observe the positions of all agents as well as doors on the map and infer the correct action based on this environmental context.
+Comparative Results of LLMs in Environment Comprehension, ToM Reasoning, and Joint Planning.
 In Figure 3, we see that most LLMs achieve their best results on the Environment Comprehension question.
 The best performing LLM GPT-4-turbo gets more than 80% Environment Comprehension Questions correct.
-The overall performance across LLMs drops on the more challenging Theory of Mind reasoning questions, but GPT-4-turbo is still competent, reaching a 54% accuracy.
+The overall performance across LLMs drops on the more challenging Theory of Mind reasoning questions.
+Both GPT-4-turbo and GPT-4o do well on the Theory of Mind reasoning questions.
 The overall accuracy of LLMs on Joint Planning questions is still significantly weak, with even the best LLM scoring less than 40%, indicating a large room for improvement in LLMs’ ability to perform coordination reasoning.
-Another cause for concern is that open-source LLMs perform abysmally at Joint Planning, with some models performing worse than a random baseline.
-Variables | ||
-ToM | 0.813 | 0.389 |
-EC | 0.895 | 0.506 |
-Impact of Environment Comprehension and ToM Reasoning abilities on Joint Planning.
-Having defined Joint Planning as the capacity of an agent to select the appropriate subsequent action based on available information, we argue that proficiency in Environment Comprehension and Theory of Mind Reasoning is crucial for adept Joint Planning, and LLMs that do well at these two will do well at JP.
-Correlation analysis of ToM and EC with JP across the data from Figure 6 reveals that ToM has a moderate positive correlation to JP, whereas EC shows a strong positive correlation.
-5 Related Work
-Multi-agent Coordination
-In Game Theory, Pure Coordination games are situations where the payoff is commonly shared between all agents.
-In such situations, cooperating is the best strategy.
-Various benchmarks and games have been used to evaluate Multi-Agent Coordination abilities over the years including Multiparticle Environment Lowe et al.
-(2017), Overcooked-AI Carroll et al.
-(2019a), and the Hanabi Challenge Bard et al.
-(2020).
-The foundational work by Carroll et al.
-(2019a) emphasized the significance of incorporating human data for effective human-ai collaboration.
-Subsequent research on the Overcooked-AI challenge has pivoted towards enabling self-play-trained agents to coordinate seamlessly with humans within this environment.
-These studies employ various techniques, including self-play with past agent checkpoints (Strouse et al., 2021), centralized population entropy objectives (Zhao et al., 2023), open-ended objectives using graph theory (Li et al., 2023a), policy ensembles with context-aware mechanisms (Lou et al., 2023), and the incorporation of human biases as linear hidden rewards (Yu et al., 2023), to enhance the training and diversity of AI agents in different scenarios.
-On the Hanabi Challenge much effort has been made to learn grounded policies Hu et al.
-(2021b; a) over arbitrary conventions.
-Embodied environments usually set up in household environments have also been recently used to study multi-agent coordination (Puig et al., 2021; Jain et al., 2020; 2019; Gan et al., 2021).
-The Overwhelming majority of approaches to coordination problems have focused on utilizing and enhancing Reinforcement Learning methods to solve the problems of multi-agent coordination.
-In this work, we argue that Large Language Models are an alternative approach to these coordination problems as they show emergent reasoning abilities, demonstrate theory-of-mind-like abilities, and do not converge to policies that cause arbitrary joint interactions.
-Planning and Reasoning with Large Language Models
-Large Language Models (LLMs) have demonstrated remarkable capabilities of reasoning in natural language (OpenAI, 2023; Ouyang et al., 2022; Chiang et al., 2023), achieving state-of-the-art performance across a spectrum of verbal reasoning tasks.
-It was then shown that LLMs could be augmented with components like memory, tools, perception, and grounding to create agents that could interact with an external environment (the web, simulators, games, etc.) These LLM agents have shown to be capable of solving long-horizon tasks, playing complex games (Wu et al., 2023; Wang et al., 2023) and interacting with simulated embodied environments (Liang et al., 2022; Song et al., 2022).
-Zhang et al.
-(2023b) developed a modular agent framework that was capable of cooperating with partner agents in embodied spatial rearrangement problems, demonstrating increased efficiency through coordination.
-Zhang et al.
-(2023a) develop a specialized architecture that enables LLMs to play in the Overcooked-AI environment.
-Li et al.
-(2023b) evaluate and show emergent collaborative abilities of LLMs in gamified simulations.
-In contrast to existing works, our work focuses on evaluating language agents in established pure coordination games where coordination is not an optional efficiency enhancer but rather a necessity.
+Another cause for concern is that open-source LLMs perform abysmally at Joint Planning, with some models performing worse than random.
+LLM Agents are robust to unseen partners.
+We use Overcooked-AI and the Hanabi challenge as testbeds to evaluate the performance of LLM agents when paired with unseen agents.
+This task is popularly known as Zero Shot Coordination.
+In Overcooked-AI, we pair our LLM agents as well as baselines with proxy-human agents.
+These proxy human agents are behavior cloning agents trained using human data by Carroll et al.
+(2019b).
+As shown in Table 4, we discover that LLM agents outperform both Behavior Cloning as well as PPO agents trained with human data.
+Apart from the Asymmetric Advantages layout, they also match or outperform the Hidden Utility Self-play (HSP) baseline, which is designed to excel at ZSC.
+In Hanabi, we pair our agents with Off-Belief Learning (OBL) agents (Hu et al., 2021a).
+OBL is a MARL strategy that generates grounded clues and actions and is the state-of-the-art method for cross-play in Hanabi.
+OBL agents provide observation-grounded clues and collaborate well with humans.
+Therefore, we use them as unseen partners in our experiments.
+Table 5 shows that the GPT-4-turbo agent scores an average of 15 points with the OBL-1 agent compared to their self-play scores of 13.66, indicating no degradation in performance with a new partner.
+The baseline RL method, Simplified Action Decoder (SAD) Hu and Foerster (2021), fails critically when paired with unseen OBL agents, even though it excels at self-play (22.00 points) due to self-play training.
 6 Conclusion
-In this study, we evaluated and analyzed the current large language models in terms of their ability to reason and act in pure coordination games.
+In this study, we evaluated and analyzed the current large language models in the context of pure coordination games.
 We introduced the LLM-Coordination benchmark with its two tasks: 1.
@@
 CoordinationQA.
-These settings allowed us to conduct holistic comparative studies of LLMs as agents as well as dive deeper into the fine-grained aspects of LLMs as coordination reasoners.
+These settings allowed us to conduct holistic comparative studies of LLMs as agents and dive deeper into the fine-grained aspects of LLMs as coordination reasoners.
 We juxtaposed LLM agents with existing Multi-agent Reinforcement Learning agents, discussing the conditions in which LLMs thrive and fail.
-Finally, we discussed the Theory of Mind Reasoning and Environment Comprehension as prerequisites for coordination and evaluated existing LLMs on these two components.
+Finally, we discussed the Theory of Mind Reasoning, Environment Comprehension, and Joint Planning as prerequisites for coordination and evaluated existing LLMs on these components.
+7 Acknowledgements
+This research project has benefitted from the Microsoft Accelerate Foundation Models Research (AFMR) grant program.
+8 Limitations
+Latency and Compute Requirements: As highlighted in Section 5, effective reasoning for coordination is achievable primarily with larger LLMs like GPT-4-turbo.
+However, these models are associated with significant latency and require substantial computational resources, making them less suitable for real-time applications where rapid decision-making is crucial.
+Initial Prompt Configuration: Achieving optimal reasoning performance necessitates careful manual configuration of the initial prompts that describe the game (Procedural Memory).
+While this prompt could be extracted from game manuals or existing resources, it still needs to be formatted and designed with the LLM agent in mind.
+Furthermore, the results for individual games could be improved by letting the LLM generate more text and engineering the prompt.
+However, we leave these optimizations to future works focused on performance improvement rather than benchmarking.
+Manual Curation of Edge Cases: The CoordinationQA suite involves manually curating unambiguous edge cases in coordination games to construct the dataset.
+This can hinder the ability to scale the benchmark to accommodate new scenarios.
+Yet, it is important to curate these examples for more reliable studies.
 Appendix A Overcooked Implementation Details
+In the game Overcooked-AI, two chefs must coordinate their actions to cook and deliver onion soups.
+Each soup requires three onions as ingredients, which can be found in onion dispensers.
+The players must add three onions to a cooker to start cooking.
+Once the three onions are added, the soup starts cooking automatically and requires 20 time steps to complete.
+Once the soup is cooked, a player needs to pick up a plate, load the soup, and deliver it to the required delivery area.
+Players must coordinate their actions to effectively deliver soup.
+Depending on the layouts (see figure 4), players need to adapt to cramped spaces, understand environment layouts for role assignments, pass objects to each other, make way for one another, and find the most effective paths.
 A.1 Game and Layout Description
@@
 The state description begins by detailing the inventories and of Alice and Bob, respectively.
-Each inventory (where ) can contain one of the following items: {”onion”, ”plate”, ”cooked soup”}.
+Each inventory (where ) can contain one of the following items: {"onion", "plate", "cooked soup"}.
 This inventory information is translated into natural language and incorporated into in the format: “I am holding .
@@
 Specifically, for each cooker , includes the number of onions currently in the pot.
-Additionally, provides the operational state of the cooker, denoted as , which can be either ”Off” or ”On”.
-Lastly, the current condition of the soup in the cooker is represented by , which can take one of the following values: ”Cooking”, ”Cooked”, or ”Not Started”.
+Additionally, provides the operational state of the cooker, denoted as , which can be either "Off" or "On".
+Lastly, the current condition of the soup in the cooker is represented by , which can take one of the following values: "Cooking", "Cooked", or "Not Started".
 Thus, the information for cooker is formatted as: “ has onions.
@@
 Appendix B Hanabi Implementation Details
+Hanabi is a card game in which all players are on the same team.
+The deck is made up of cards numbered 1 through 5, further divided into five different colors.
+Players are working together to create these numbered sequences (1 through 5) for each of the five colors.
+Each card played provides 1 point, and the goal is to obtain all 25 points (5 colors with 5 cards each) by completing the sequence for each color.
+The three actions a player can take include playing a card, discarding a card, and giving a hint (reveal) to a teammate.
+The challenge is that players cannot see their own cards.
+They may, however, see the cards that are in the hands of their teammate(s).
+This is where hints (reveals) come into play.
+Players are able to figure out which cards should be played through the reveal mechanism.
+You can hint (reveal) a teammate about a color or number in their hand, and this will point out to them all cards in their hand that have this color or number.
+The team starts with 8 reveal tokens but can recover used tokens when discarding a card or upon the completion of a stack.
+Playing a card carries a risk because the team loses completely if three incorrect cards are played.
+Further, an incorrectly played card gets sent to the discard pile.
+Each number of a particular color only has so many copies, so if, for example, the last copy of a green 3 is lost, then the team loses out on not only playing the green 3 but also the green 4 and 5.
+This is the risk of playing or discarding.
+With these risks, and since reveal tokens are scarce, players ideally try to make assumptions about implicit information that a reveal may offer, and, ideally, the team converges to particular conventions.
+Implicit communication, collaboration, and memory are key.
 B.1 Game Description
-We structure the game description of Hanabi into the overall objective, the rules of the game, and list of conventions based on the H-group conventions han (2024).
-We do not use advanced conventions like Chop Cards but stick to basic conventions about the card layout, play clues, and save clues.
+We structure the game description of Hanabi into the overall objective, and the rules of the game.
 B.2 State Description
@@
 We also precalculate the next card that goes on each stack since LLMs frequently fail to count which card should go next on each stack.
-Appendix C Examples of prompts of LLMs used in CAC framework
-C.1 Action Generator
-The action generator generates a brief explanation and selected the next action to be played.
-C.2 Theory of Mind Reasoning LLM
-The Theory of Mind Reasoning LLM interprets partner actions and clarifies partner requirements.
-C.3 Self Verification LLM
-The prompt for the Self Verification LLM is provided as a system prompt. We observed that LLMs were more likely to act as strict verifiers when the prompt is provided as a system prompt.
-Appendix D Results of Different LLMs on CollabCapture and CollabEscape
-Table 7 summarizes the performance of the three LLMs GPT-4-turbo, GPT-3.5-turbo and Mixtral-8x7b on CollabGames - Collab Capture and Collab Escape.
-GPT-4-turbo achieves a 100% success rate in both games.
-In general, CollabEscape was a more difficult game to solve than CollabGames since agents needed to perform sacrificial moves and lure the killer away, requiring more advanced Theory of Mind inference.
-Collab Capture | Collab Escape | |||
-CAC model | Capture Rate | Turns to Capture | Escape Rate | Turns to Escape |
-GPT-4-turbo | 1.00 | 3.99 | 1.00 | 24.67 |
-GPT-3.5-turbo | 0.50 | 8.49 | 0.00 | N.A. |
-Mixtral-8x7b | 0.75 | 3.88 | 0.00 | N.A. |
-Greedy Baseline | 0.50 | 6.00 | 0.00 | N.A. |
-Appendix E Generating Questions for CoordinationQA
-E.1 Environment Comprehension Questions
+Appendix C Examples of prompts of LLMs used in the Agent framework
+C.1 Hanabi LLM Prompt
+C.2 Prompt for Theory of Mind Reasoning step
+C.3 Prompt for Answer Verification Step
+C.4 Ablation on Theory of Mind Reasoning and Verification Steps
+| Method | Score | Bomb Rate |
+| ToM + GPT-4-turbo + V | ||
+| GPT-4-turbo + V | ||
+| GPT-4-turbo |
+Appendix D Generating Questions for CoordinationQA
+D.1 Environment Comprehension Questions
 The Environment Comprehension (EC) questions are indirect formulations regarding spatial aspects of the layout.
-In order for an agent to correctly answer an EC question, they must have an understanding of the dynamic details of the current state, the rules of the game, and exhibit spatial awareness.
+In order for an agent to correctly answer an EC question, they must have an understanding of the dynamic details of the current state and the rules of the game and exhibit spatial awareness.
 As such, when creating the EC questions, we carefully comb through a given scenario in search of salient points to probe an agent’s understanding of the given environment.
 Some examples include:
-E.2 Theory of Mind Reasoning Questions
+D.2 Theory of Mind Reasoning Questions
 There are two primary question types in Hanabi for ToM Reasoning questions.
@@
 For all games apart from Hanabi, the ToM questions ask the models to predict the next intended action of the partner agent
-E.2.1 Hanabi Question Type-1
-E.2.2 Hanabi Question Type-2
-E.2.3 Other Games
-E.3 Joint Planning Questions
+D.2.1 Hanabi Question Type-1
+D.2.2 Hanabi Question Type-2
+D.2.3 Other Games
+D.3 Joint Planning Questions
 Joint planning questions are effectively the same questions that the LLM solves when they are part of an agentic framework.
-For each scenario, we ask the LLM to answer the question: ”What is the best next action?”.
+For each scenario, we ask the LLM to answer the question: "What is the best next action?".
+Appendix E CollabCapture and CollabEscape
+E.1 CollabCapture
+CollabCapture places two agents in a set of interconnected rooms and doors.
+Their goal is to capture an adversary within the smallest number of moves possible.
+The adversary moves away from agents in a greedy fashion, but the agents have the ability to close and open doors while the adversary does not.
+The doors are controlled by a corresponding button that is in another location on the map.
+(See Figure 5)
+CollabCapture contains one layout with four scenarios created by controlling the gate states (open/closed).
+This corresponds to cases where players need to: 1.
+Pincer their opponent through coordination 2.
+One agent needs to enable the other agent by choosing not to chase the agent but rather open the door to allow the other agent to capture the adversary.
+3.
+One agent needs to disable the adversary by closing a door, allowing the other agent to catch them.
+CollabCapture is based on the classic task of Pursuit Evasion from the perspective of the pursuers.
+This is representative of a common-payoff task as the only objective is capturing the adversary with no mixed incentives (Akin to deliveries in Overcooked, where all chefs get a common payoff with no preference for the one making the delivery).
+E.2 CollabEscape
+CollabEscape also places two agents in a set of interconnected rooms and doors.
+Their goal in this environment, however, is to escape an adversary that is looking to catch them.
+The map has two generators that power an exit gate, both of which need to be fixed.
+Upon fixing both generators, the players must then escape by reaching the exit gate.
+Only one player needs to reach the exit gate in order for them both to win.
+If the adversary catches either of them, however, they both lose.
+(See Figure 6)
+In CollabEscape, we have one layout with two scenarios created by varying the starting positions of the two agents.
+Depending on their proximity to the adversary/generators, players need to apply strategies such as luring the adversary away from the partner, choosing to continue fixing the generators while sacrificing for the partner’s safety, and manipulating the movement of the adversary.
+This is also representative of a common payoff task, as players are commonly rewarded if any one of them escapes, introducing roles of explicit assistance and sacrifice for a higher common payoff.